OpenStudy (anonymous):

need help on this problem... 5= -3x - 3y -3y = 3x - 2 these are my choices ... no solution infinite solutions (0,6) (5,-3)

OpenStudy (anonymous):

this is solving simultaneous equations, have u done that before?

OpenStudy (anonymous):

not really??......

OpenStudy (anonymous):

try read this first. http://www.mathsrevision.net/gcse-maths-revision/algebra/simultaneous-equations

OpenStudy (anonymous):

then try do solve it urself and check with my working out that I'm going to provide.

OpenStudy (anonymous):

too much reading lol

OpenStudy (anonymous):

READ IT!!! lol

OpenStudy (anonymous):

ahhh lol fine

OpenStudy (anonymous):

which choice would you pick????

OpenStudy (anonymous):

u have to read the link provided and then solve it urself first.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

to figure out the answer.

OpenStudy (anonymous):

5 = -3x - 3y 3y = -3x + 5 y = -x + 5/3 -3y = 3x - 2 y = -x + 2/3 the two equations have the same gradient or m indicated by the coefficient in front of x but the y-intercept is different so therefore the answer will be no solution.